# Unsupervised Interpretable Basis Extraction for Concept – Based Visual Explanations

### v2

@@ 6.1 Basis labeling and classifier validation scores
To quantitatively evaluate for interpretability the bases extracted with our method, we use a two step process. First, after deriving a basis, we use the work of Bau at al. [11, 28], to assign a concept label to each classifier associated with the basis vectors. Let $\phi(i,c,\mathcal{K})\in[0,1]$ denote a metric score function that is used to measure the suitability of the classifier $i$ ($\{\mathbf{w}_{i},b_{i}\}$) to accurately detect concept $c$ in the annotated concept dataset $\mathcal{K}$. The concept label that is assigned to classifier $i$ is the one that maximizes $\phi(i,c,\mathcal{K}_{train})$ across $c$ over the training split $\mathcal{K}_{train}$ of the concept dataset. Subsequently, in the second step, and using the validation split of the concept dataset $\mathcal{K}_{val}$, each classifier is assigned a validation score $\phi(i,c^{\star}_{i},\mathcal{K}_{val})$, with $c^{\star}_{i}$ denoting the concept label assigned to the classifier during the first step. For the choice of $\phi$ we use Intersection Over Union (IoU), as originally proposed in [11] and also used in [12, 9]:
@@
In (8), $\mathbf{M}^{i}(\mathbf{k})$ denotes the upsampled, hard-thresholded (binarized) map of image $\mathbf{k}$. $\mathbf{M}^{i}(\mathbf{k})$ is obtained by applying the rule of the $i$-th classifier ($\mathbf{w}_{i}^{T}\mathbf{x_{p}}-b_{i}>0)$ to each $\mathbf{x_{p}}$ of the upsampled image’s representation. Moreover, $\mathbf{L}^{c}(\mathbf{k})$ denotes the ground truth segmentation map of image $\mathbf{k}$ for concept $c$ and ||$\cdot$|| ⋅ | denotes the cardinality of a set. Overall, to label the bases and compute classifier validation scores, we use the exact scheme of [11] with two differences. First, we consider a train/test split of the concept dataset as originally proposed in [9] and second, for hard-thresholding in $\mathbf{M}^{i}(\mathbf{k})$, we use the biases learned from our method, instead of using the statistical quantile learning of [11].
@@ 7 Experimental Results
Basis Learning Details To learn each one of the basis, we used the Adam [29] optimizer with the default beta parameters (0.9, 0.9990.90.9990.9,\,0.9990.9 , 0.999) provided by the PyTorch [30] implementation. We fixed the learning rate to $0.001$ and did not employ any form of learning rate scheduling. In all cases, basis learning lasted for $300$ epochs. Batch size was a variable that varied across our experiments and its value was based solely on the available GPU memory resources. The values we used, approximately lied in the interval $\approx[800-3600]$.
@@ 7.2 Results for more networks
Figure 13: Ablation with respect to the number of partitions $N$. For ResNet18, just a single partition resulted into the most interpretable basis. However, for other values of $N$ the results are comparable and an improvement is noted compared to the baseline.
Figure 14: Ablation with respect to the number of partitions $N$. For VGG16BN, four partitions resulted into the most interpretable basis. However, for other values of $N$ the results are comparable and an improvement is noted compared to the baseline.
@@ 7.3 Comparison with a supervised approach
Regarding $\mathcal{S}^{2}$, the bases extracted with IBD may be considered significantly more interpretable than the bases extracted with the current work, with the latter being even more prominent in the case of ResNet18. We think that this fact is also linked to the previous argument. In particular, since a basis learned without supervision may have duplicate labels, the number of unique concept labels that can be attributed to the basis vectors (which is related to what $\mathcal{S}^{2}$ measures), is expected to be less than the number of vectors in the basis. However, for a basis learned in a supervised way, this number is exactly equal to the number of vectors in the basis ($I$). Moreover, in this case, IBD used a basis with a larger number of vectors ($I=660$) compared to the proposed method (which only uses $I=D=512$).
@@ 8 Conclusion
We presented an unsupervised, post-hoc method to extract an interpretable basis for a CNN’s intermediate layer feature space. Based on current literature, we also proposed two metrics that can be used to measure a basis for its interpretability. We evaluated the effectiveness of the proposed method in standard CNN architectures and demonstrated that intermediate layer representations become more interpretable when projected onto bases extracted with our method. Finally, using the proposed metrics, we compared the outcomes of our method with the outcomes of a method that derives an interpretable basis using supervision. According to the interpretability metrics, the bases extracted with the proposed method, in one aspect, show appreciable interpretability improvements over the bases extracted with the supervised approach. At the same time, in a second aspect, the bases derived with supervision, were significantly more interpretable than the bases that were suggested by our method. This fact might seem peculiar at first. However, a possible explanation was provided and directions for future research for deeper understanding were suggested. We hope that the present work has contributed additional knowledge to interpretable basis extraction and motivates further research for understanding black-box models.

### v3

@@ 6.1 Basis labeling and classifier validation scores
To quantitatively evaluate the bases extracted with our method for interpretability, we use a two step process. First, after deriving a basis, we use the work of Bau at al. [11, 28], to assign a concept label to each classifier associated with the basis vectors. Let $\phi(i,c,\mathcal{K})\in[0,1]$ denote a metric score function that is used to measure the suitability of the classifier $i$ ($\{\mathbf{w}_{i},b_{i}\}$) to accurately detect concept $c$ in the annotated concept dataset $\mathcal{K}$. The concept label that is assigned to classifier $i$ is the one that maximizes $\phi(i,c,\mathcal{K}_{train})$ across $c$ over the training split $\mathcal{K}_{train}$ of the concept dataset. Subsequently, in the second step, and using the validation split of the concept dataset $\mathcal{K}_{val}$, each classifier is assigned a validation score $\phi(i,c^{\star}_{i},\mathcal{K}_{val})$, with $c^{\star}_{i}$ denoting the concept label assigned to the classifier during the first step. For the choice of $\phi$ we use Intersection Over Union (IoU), as originally proposed in [11] and also used in [12, 9]:
@@
In (8), $\mathbf{M}^{i}(\mathbf{k})$ denotes the upsampled, hard-thresholded (binarized) map of image $\mathbf{k}$. $\mathbf{M}^{i}(\mathbf{k})$ is obtained by applying the rule of the $i$-th classifier ($\mathbf{w}_{i}^{T}\mathbf{x_{p}}-b_{i}>0$) to each $\mathbf{x_{p}}$ of the image’s representation followed by upsampling to the resolution of the original image. Moreover, $\mathbf{L}^{c}(\mathbf{k})$ denotes the ground truth segmentation map of image $\mathbf{k}$ for concept $c$ and $|\cdot|$ denotes the cardinality of a set. Overall, to label the bases and compute classifier validation scores, we use the exact scheme of [11] with two differences. First, we consider a train/test split of the concept dataset as originally proposed in [9] and second, for hard-thresholding in $\mathbf{M}^{i}(\mathbf{k})$, we use the biases learned from our method, instead of using the statistical quantile learning of [11].
@@ 7 Experimental Results
Basis Learning Details To learn each one of the basis, we used the Adam [29] optimizer with the default beta parameters (0.9, 0.9990.9,\,0.999) provided by the PyTorch [30] implementation. We fixed the learning rate to $0.001$ and did not employ any form of learning rate scheduling. In all cases, basis learning lasted for $300$ epochs. Batch size was a variable that varied across our experiments and its value was based solely on the available GPU memory resources. The values we used, approximately lied in the interval $\approx[800-3600]$.
@@ 7.2 Results for more networks
Figure 13: Ablation with respect to the number of partitions NN. For ResNet18, just a single partition resulted into the most interpretable basis. However, for other values of $N$ the results are comparable and an improvement is noted compared to the baseline.
Figure 14: Ablation with respect to the number of partitions NN. For VGG16BN, four partitions resulted into the most interpretable basis. However, for other values of $N$ the results are comparable and an improvement is noted compared to the baseline.
@@ 7.3 Comparison with a supervised approach
Regarding $\mathcal{S}^{2}$, the bases extracted with IBD may be considered significantly more interpretable than the bases extracted with the current work, with the latter being even more prominent in the case of ResNet18. We think that this fact is also linked to the previous argument. In particular, since a basis learned without supervision may have duplicate labels, the number of unique concept labels that can be attributed to the basis vectors (which is related to what $\mathcal{S}^{2}$ measures) is expected to be less than the number of vectors in the basis. However, for a basis learned in a supervised way, these two numbers are always equal. Moreover, in this case, IBD used a basis with a larger number of vectors ($I=660$) compared to the proposed method (which only uses $I=D=512$).
@@ 8 Conclusion
We presented an unsupervised, post-hoc method to extract an interpretable basis for the feature space of a CNN’s intermediate layer. Based on current literature, we also proposed two metrics that can be used to measure a basis for its interpretability. We evaluated the effectiveness of the proposed method in standard CNN architectures and demonstrated that intermediate layer representations become more interpretable when projected onto bases extracted with our method. Finally, using the proposed metrics, we compared the outcomes of our method with the outcomes of a method that derives an interpretable basis using supervision. According to the interpretability metrics, the bases extracted with the proposed method, in one aspect, show appreciable interpretability improvements over the bases extracted with the supervised approach. At the same time, in a second aspect, the bases derived with supervision were significantly more interpretable than the bases that were suggested by our method. This fact might seem peculiar at first. However, a possible explanation was provided and directions for future research for deeper understanding were suggested. We hope that the present work has contributed additional knowledge to interpretable basis extraction and motivates further research for understanding black-box models.